OpenStudy (anonymous):

Whats is 100+(4^5-6^2)=

OpenStudy (solomonzelman):

4^5=1024 6^2=36 So, (4^5-6^2) = (1024-36) = 988 Now, 100+988 = 1088 there is your answer, see how I did it ?